OpenStudy (anonymous):

what is the theoretical probability of being dealt all four 7's ina 5 card hand?

jimthompson5910 (jim_thompson5910):

Are you familiar with the combination formula?

OpenStudy (dumbcow):

probability = (possible hands with four 7's)/ (all possible hands) total num of hands = 52 choose 5 num of hands w/ four 7's = (4 choose 4)*(48 choose 1) \[P = \frac{\left(\begin{matrix}4 \\ 4\end{matrix}\right)*\left(\begin{matrix}48 \\ 1\end{matrix}\right)}{\left(\begin{matrix}52 \\ 5\end{matrix}\right)}\] formula for combinations \[\left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{n!}{k! (n-k)!}\]